# Faces of homogeneous cones and applications to homogeneous chordality

### v2

@@ 4 Applications to homogeneous chordality
and we denote by $\mathcal{S}(G)$ the space spanned by $\mathcal{S}_{+}({G})$. If $G$ is chordal, the cone $\mathcal{S}_{+}({G})$ is generated by its rank one matrices, e.g., [AHMR88, Theorem 2.3]. When $G$ not only is chordal but also does not contain induced subgraphs isomorphic to a path on $4$ vertices, by [Ish13, Theorem A] they are also homogeneous and we will refer to such a graph $G$ as being a homogeneous chordal graph. This class of cones was also extensively studied in [TV23].
@@ 2. $({o}2)$
Axiom 4 needs a little more thought. We have to check that for matrices $a,b,c\in\mathcal{A}(G)$ we have $\textup{tr}((a\star b)\star c)=\textup{tr}(a\star(b\star c))$. As $\pi_{G}$ never zeroes diagonal entries, this is equivalent to checking that $\textup{tr}(\pi_{G}(ab)c)=\textup{tr}(a\pi_{G}(bc))$ holds. First we prove $\textup{tr}(\pi_{G}(ab)c)=\textup{tr}(abc)$ by showing that $\pi_{G}(ab)c$ and $abc$ have the same diagonal entries. Note that if $[\pi_{G}(ab)]_{ij}$ is different from $[ab]_{ij}$, this means $\{i,j\}$ is not in $E$, so $c_{ji}=0$. Therefore
$[\pi_{G}(ab)c]_{ii}=\sum_{j=1}^{n}[\pi_{G}(ab)]_{ij}c_{ji}=\sum_{j=1}^{n}[ab]_{ij}c_{ji}=[abc]_{ii}.$
@@
This tells us that ${\mathcal{K}}(\mathcal{A}(G))^{*}=\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}\subseteq\mathcal{S}_{+}({G})$. The converse and a summary of the discussion so far are given in the next theorem.
@@ 3. $(iii)$
It remains to verify the inclusion $\mathcal{S}_{+}({G})\subseteq\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}$ and items ii and iii. We start with the former. The extreme rays of $\mathcal{S}_{+}({G})$ are generated by rank one matrices, e.g., [AHMR88, Theorem 2.3]. So, let $v\in M^{n\times 1}$ (i.e., a column vector) be such that $vv^{*}$ generates an extreme ray of $\mathcal{S}_{+}({G})$. Letting $J\coloneqq\{i\mid v_{i}\neq 0\}$, we observe that for every pair $i\neq j$, $\{i,j\}\subseteq J$, we must have $\{i,j\}\in E$, since $(vv^{*})_{ij}\neq 0$. Let $i$ be the smallest element of $J$. Replacing $v$ with $-v$ if necessary, we may assume that $v_{i}$ is positive. Then, if we let $l$ be the lower triangular matrix that has $v$ as it is $i$-column and is zero elsewhere, we have $l\in\mathcal{L}_{+}(G)$, since $l_{ji}=v_{j}$ and $\{j,i\}\in E$ whenever $v_{j}\neq 0$ and $i\neq j$. We also have $ll^{*}=vv^{*}$ which shows that $\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}$ contains all the extreme rays of $\mathcal{S}_{+}({G})$ and leads to the required inclusion.
@@
Principal subalgebras of $\mathcal{A}(G)$ are obtained by zeroing out some rows and columns of $\mathcal{A}(G)$. This corresponds to removing vertices of the underlying homogeneous chordal graph, an operation that in fact preserves homogeneous chordality. For $H=(\hat{V},\hat{E})$ an induced subgraph of $G$, we define
@@
and, recalling Section 3.3, we observe that this coincides with the principal dual face $\mathrm{cl}\,({\mathcal{K}}^{D}(\mathcal{A}(G)_{I}))$ induced by $I=\{i\mid i\not\in\hat{V}\}$. Also, ${\mathcal{F}}(H)$ (which is a cone of $n\times n$ matrices) is linearly isomorphic to $\mathcal{S}_{+}({H})$ (which is a cone of $|\hat{V}|\times|\hat{V}|$ matrices). Theorems 3.6 and 3.10 then imply that every face of a homogeneous chordal cone $\mathcal{S}_{+}({G})$ is isomorphic to a homogeneous chordal cone $\mathcal{S}_{+}({H})$, where $H$ is an induced subgraph of $G$. In more details, we have the following theorem.
@@ 2. $(ii)$
${\mathcal{F}}$ =$=\{x\in\mathcal{S}_{+}({G})\mid(Q_{l}(x))_{33}=0\}=\{x\in\mathcal{S}_{+}({G})\mid(lxl^{*})_{33}=0\}$ (4.4)
=$=\{x\in\mathcal{S}_{+}({G})\mid x_{11}-2x_{31}+x_{22}-2x_{32}+x_{33}=0\}.$
@@ 4.1 The PSD completion side
In the next theorem, we gather several facts about the facial structure of $\pi_{G}(\mathcal{S}_{+}^{n})$ together with a proof that Algorithm 1 leads to a maximum rank decomposition. As before, if $H=(\hat{V},\hat{E})$ is an induced subgraph of $G$, then it also corresponds to a face of $\pi_{G}(\mathcal{S}_{+}^{n})$ in a natural way. By an abuse of notation, we define $\pi_{H}(\mathcal{S}_{+}^{n})\coloneqq\{x\in\pi_{G}(\mathcal{S}_{+}^{n})\mid x_{ij}=x_{ji}=0,\forall i,j\not\in\hat{V}\}$. Put otherwise, $x\in\pi_{H}(\mathcal{S}_{+}^{n})$ if it has a PSD completion $w$ satisfying $w_{ii}=0$ for $i\not\in\hat{V}$.
@@ 2. $(iv)$
For item $(iii)$, we start by observing that ${\mathcal{F}}$ is a homogeneous cone of rank $r$, also by item $(ii)$ of Theorem 3.6. Then, ${\mathcal{F}}^{\Delta}$ is a homogeneous cone of rank $n-r$, by Proposition 3.11. Let $y\in\mathrm{ri}\,{\mathcal{F}}^{\Delta}$ and write $y=ll^{*}$, where $l$ is a proper lower triangular matrix in $\mathcal{L}_{+}(G)$. By Theorem 3.10, since ${\mathcal{F}}^{\Delta}$ has rank $n-r$, the diagonal of $l$ has exactly $n-r$ positive elements, so $y$ has rank $n-r$. Let $w$ be a PSD completion of $x$, then
@@
Since $x$ generates an extreme ray, ${\mathcal{F}}^{\Delta}$ is a maximum proper face of $\mathcal{S}_{+}({G})$ by Lemma 2.1. Let $y^{\prime}\in\mathcal{S}_{+}({G})$ be such that it generates an extreme ray that is not in ${\mathcal{F}}^{\Delta}$. Since ${\mathcal{F}}^{\Delta}$ is maximal, we have $\mathrm{Face}(y+y^{\prime},\mathcal{S}_{+}({G}))=\mathcal{S}_{+}({G})$, which implies that $y+y^{\prime}$ is positive definite and, in particular, has rank $n$. Because $G$ is chordal, extreme rays correspond to rank one matrices ([AHMR88, Theorem 2.3]), so $y^{\prime}$ has rank one, which implies that $y$ has rank $n-1$.
@@
Beyond Proposition 4.4, we do not know if items $(i)$ and $(ii)$ of Theorem 4.3 have analogues for the case where $G$ is only chordal.
@@ 2. $(ii)$
    $\pi_{G}(\mathcal{S}_{+}^{n})=\mathcal{S}_{+}({G})$ (i.e., $\mathcal{S}_{+}({G})$ is self-dual)
@@ 5 Conclusion and open questions
Here we briefly describe a potential practical application. When solving a semidefinite program that fails Slater’s condition, one of the default ways of restoring constraint qualifications and reducing the size of the problem is through facial reduction [BW81, Pat13b, WM13]. This is done by reformulating the problem over a smaller face of the cone, which leads to a smaller SDP. The reformulation typically uses the projectional exposedness of $\mathcal{S}_{+}^{n}$ and expressions such as (1.1). Unfortunately, this often destroys sparsity if $q$ is as in (1.1), since $qxq^{*}$ does not necessarily satisfy any sparsity pattern present in $x$. Therefore, one must face the inevitable choice of either keeping sparsity or regularizing the problem. From a theoretical point of view, the issue is that the map $x\mapsto qxq^{*}$ is an automorphism of $\mathcal{S}_{+}^{n}$ but not necessarily an automorphism of the PSD cone that corresponds to the sparsity pattern of the problem.
@@ 1. $(a)$
    Suppose that $G=(V,E)$ is chordal but not necessarily homogeneous. Are $\mathcal{S}_{+}({G})$ and $\pi_{G}(\mathcal{S}_{+}^{n})$ projectionally exposed? Do analogues of Theorem 4.2 and Theorem 4.3 hold? For example, it would be interesting to check whether, assuming only chordality, the faces of $\mathcal{S}_{+}({G})$ are linearly isomorphic to the finitely many faces that arise by considering induced subgraphs of $G$ and if such an isomorphism can be realized through an automorphism of $\mathcal{S}_{+}({G})$. The smallest example for which we do not know the answer to these questions is when $G$ is a path on four vertices. Analogously to the previous discussion, a positive answer would have interesting implications for facial reduction approaches.

### v3

@@ 4 Applications to homogeneous chordality
and we denote by $\mathcal{S}(G)$ the space spanned by $\mathcal{S}_{+}({G})$. If $G$ is chordal, the cone $\mathcal{S}_{+}({G})$ is generated by its rank one matrices, e.g., [AHMR88, Theorem 2.3]. Chordal graphs have found extensive applications in solving semidefinite programs efficiently, e.g., [FKMN01, NFF+03, VA15]. In particular, it is well-known that, by reordering rows and columns if necessary, positive definite matrices with a chordal sparsity pattern have Cholesky factorizations that respect the underlying sparsity pattern, i.e., factorizations with “zero fill-in”.
When $G$ not only is chordal but also does not contain induced subgraphs isomorphic to a path on $4$ vertices, by [Ish13, Theorem A] they are also homogeneous and we will refer to such a graph $G$ as being a homogeneous chordal graph. This class of cones was also extensively studied in [TV23]. A graph that does not contain an induced subgraph isomorphic to a path on $4$ vertices is also called a cograph, so, alternatively, a homogeneous chordal graph can be defined as a chordal cograph. Cholesky factorizations under homogeneous chordal sparsity have stronger properties when compared to plain chordal sparsity, e.g., inverses also respect the underlying sparsity pattern [TV23, item (ii) of Theorem 3.1].
@@
Axiom 4 needs a little more thought. We have to check that for matrices $a,b,c\in\mathcal{A}(G)$ we have $\textup{tr}((a\star b)\star c)=\textup{tr}(a\star(b\star c))$. As $\pi_{G}$ never zeroes diagonal entries, this is equivalent to checking that $\textup{tr}(\pi_{G}(ab)c)=\textup{tr}(a\pi_{G}(bc))$ holds. First we prove $\textup{tr}(\pi_{G}(ab)c)=\textup{tr}(abc)$ by showing that $\pi_{G}(ab)c$ and abcabc have the same diagonal entries. Note that if $\pi_{G}(ab)]_{ij}$ is different from $ab]_{ij}$, this means $\{i,j\}$ is not in $E$, so $c_{ji}=0$. Therefore
$\pi_{G}(ab)c]_{ii}=\sum_{j=1}^{n}[\pi_{G}(ab)]_{ij}c_{ji}=\sum_{j=1}^{n}[ab]_{ij}c_{ji}=[abc]_{ii}.$
@@
This tells us that ${\mathcal{K}}(\mathcal{A}(G))^{*}=\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}\subseteq\mathcal{S}_{+}({G})$. The converse and a summary of the discussion so far are given in the next theorem. We note that items $(i)$ and $(iii)$ below also follow from the [TV23, Theorem 3.1].
@@ 3. $(iii)$
It remains to verify the inclusion $\mathcal{S}_{+}({G})\subseteq\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}$ and items ii and iii. We start with the former. The extreme rays of $\mathcal{S}_{+}({G})$ are generated by rank one matrices, e.g., [AHMR88, Theorem 2.3]. So, let $v\in M^{n\times 1}$ (i.e., a column vector) be such that $vv^{*}$ generates an extreme ray of $\mathcal{S}_{+}({G})$. Letting $J\coloneqq\{i\mid v_{i}\neq 0\}$, we observe that for every pair $i\neq j$, $\{i,j\}\subseteq J$, we must have $\{i,j\}\in E$, since $(vv^{*})_{ij}\neq 0$. Let $i$ be the smallest element of $J$. Replacing $v$ with v-v if necessary, we may assume that $v_{i}$ is positive. Then, if we let $l$ be the lower triangular matrix that has $v$ as it is $i$-column and is zero elsewhere, we have $l\in\mathcal{L}_{+}(G)$, since $l_{ji}=v_{j}$ and $\{j,i\}\in E$ whenever $v_{j}\neq 0$ and $i\neq j$. We also have $ll^{*}=vv^{*}$ which shows that $\{ll^{*}\mid l\in\mathcal{L}_{+}(G)\}$ contains all the extreme rays of $\mathcal{S}_{+}({G})$ and leads to the required inclusion.
@@
[…]
@@
[…]
@@ 2. $(ii)$
We move on to an example and, after that, we will observe that a part of Theorem 4.2 may fail for chordal graphs that are not homogeneous.
@@
${\mathcal{F}}$ $=\{x\in\mathcal{S}_{+}({G})\mid(Q_{l}(x))_{33}=0\}=\{x\in\mathcal{S}_{+}({G})\mid(lxl^{*})_{33}=0\}$ (4.4)
$=\{x\in\mathcal{S}_{+}({G})\mid x_{11}-2x_{31}+x_{22}-2x_{32}+x_{33}=0\}.$
@@
Remark 4.3 (Failure of Theorem 4.2 for chordal graphs).
Let $G$ be the following graph.
11223344
The graph $G$ is chordal but not homogeneous. We consider the cone $\mathcal{S}_{+}({G})$ and we will show that it has a face linearly isomorphic to $\mathbb{R}^{3}_{+}$. Let $v\coloneqq(1,1,1,1)\in\mathbb{R}^{4}$ and consider the rank $1$ matrix $vv^{*}$, which belongs to $(\mathcal{S}_{+}({G}))^{*}$. Let ${\mathcal{F}}\coloneqq\mathcal{S}_{+}({G})\cap\{vv^{*}\}^{\perp}=\{x\in\mathcal{S}_{+}({G})\mid v\in\ker x\}$.
Every matrix that generates an extreme ray of $\mathcal{S}_{+}({G})$ must have its support contained in a maximal clique of $G$, e.g., see [AHMR88, Theorem 2.3]. So if $x$ is an extreme ray of $\mathcal{S}_{+}({G})$ then it has one of the following patterns of nonzero entries:
$\begin{pmatrix}x_{11}&x_{12}&0&0\\ x_{12}&x_{22}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},\begin{pmatrix}0&0&0&0\\ 0&x_{22}&x_{23}&0\\ 0&x_{23}&x_{33}&0\\ 0&0&0&0\end{pmatrix},\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&x_{33}&x_{34}\\ 0&0&x_{34}&x_{44}\end{pmatrix}$.
If, in addition, $x$ is an extreme ray of $\mathcal{F}$} then $v$ is in the kernel of $x$ which implies that $x$ must be a positive multiple of one of the following matrices
$\begin{pmatrix}1&-1&0&0\\ -1&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},\begin{pmatrix}0&0&0&0\\ 0&1&-1&0\\ 0&-1&1&0\\ 0&0&0&0\end{pmatrix},\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&-1\\ 0&0&-1&1\end{pmatrix}$.
Each of these three matrices is also in $\mathcal{F}$}, so we conclude that $\mathcal{F}$} has exactly three extreme rays and, is therefore, linearly isomorphic to $\mathbb{R}^{3}_{+}$.
The face $\mathcal{F}$} also contains a rank $3$ matrix (e.g., the sum of these three matrices). In order for a face associated to an induced subgraph of $G$ to contain a rank $3$ matrix it must be the result of removing a single vertex. However, removing a single vertex of $G$ leads to a face that will, at the very least, contain a face isomorphic to $\mathcal{S}_{+}^{2}$ (corresponding to one of the remaining edges of the induced subgraph). In particular, such a face is not polyhedral.
The conclusion is that no face associated to an induced subgraph of $G$ can be linearly isomorphic to $\mathcal{F}$}. In particular, $Q$ as in Theorem 4.2 may not exist when $G$ is only chordal.
@@ 4.1 The PSD completion side
[…]
@@ 2. $(iv)$
For item $(iii)$, we start by observing that $\mathcal{F}$} is a homogeneous cone of rank $r$, also by item $(ii)$ of Theorem 3.6. Then, ${\mathcal{F}}^{\Delta}$ is a homogeneous cone of rank nrn-r, by Proposition 3.12. Let $y\in\mathrm{ri}\,{\mathcal{F}}^{\Delta}$ and write $y=ll^{*}$, where $l$ is a proper lower triangular matrix in $\mathcal{L}_{+}(G)$. By Theorem 3.11, since ${\mathcal{F}}^{\Delta}$ has rank nrn-r, the diagonal of $l$ has exactly nrn-r positive elements, so $y$ has rank nrn-r. Let $w$ be a PSD completion of $x$, then
@@
Since $x$ generates an extreme ray, ${\mathcal{F}}^{\Delta}$ is a maximum proper face of $\mathcal{S}_{+}({G})$ by Lemma 2.1. Let $y^{\prime}\in\mathcal{S}_{+}({G})$ be such that it generates an extreme ray that is not in ${\mathcal{F}}^{\Delta}$. Since ${\mathcal{F}}^{\Delta}$ is maximal, we have $\mathrm{Face}(y+y^{\prime},\mathcal{S}_{+}({G}))=\mathcal{S}_{+}({G})$, which implies that $y+y^{\prime}$ is positive definite and, in particular, has rank $n$. Because $G$ is chordal, extreme rays correspond to rank one matrices ([AHMR88, Theorem 2.3]), so $y^{\prime}$ has rank one, which implies that $y$ has rank n1n-1.
@@
Beyond Proposition 4.5, we do not know if items $(i)$ and $(ii)$ of Theorem 4.4 have analogues for the case where $G$ is only chordal, although Remark 4.3 may be an indication that an analogue may not exist.
@@ 2. $(ii)$
    $\pi_{G}(\mathcal{S}_{+}^{n})=\mathcal{S}_{+}({G})$ (i.e., $\mathcal{S}_{+}({G})$ is self-dual).
@@ 5 Conclusion and open questions
Here we briefly describe a potential practical application. When solving a semidefinite program that fails Slater’s condition, one of the default ways of restoring constraint qualifications and reducing the size of the problem is through facial reduction [BW81, Pat13b, WM13]. This is done by reformulating the problem over a smaller face of the cone, which leads to a smaller SDP. The reformulation typically uses the projectional exposedness of $\mathcal{S}_{+}^{n}$ and expressions such as (1.1). Unfortunately, this often destroys sparsity if $q$ is as in (1.1), since $qxq^{*}$ does not necessarily satisfy any sparsity pattern present in $x$. Therefore, one must face the inevitable choice of either keeping sparsity or regularizing the problem. From a theoretical point of view, the issue is that the map $x\mapsto qxq^{*}$ is an automorphism of $\mathcal{S}_{+}^{n}$ but not necessarily an automorphism of the PSD slice that corresponds to the sparsity pattern of the problem.
@@ 1. $(a)$
    Suppose that $G=(V,E)$ is chordal but not necessarily homogeneous. Are $\mathcal{S}_{+}({G})$ and $\pi_{G}(\mathcal{S}_{+}^{n})$ projectionally exposed? What are the connections between the induced subgraphs of $G$ and the faces of $\mathcal{S}_{+}({G})$? We already saw in Remark 4.3 that, in contrast to homogeneous chordal graphs, $\mathcal{S}_{+}({G})$ may have faces that are not linearly isomorphic to faces arising from induced subgraphs. Nevertheless, it would be interesting to clarify the extent to which the graph structure of $G$ influences the facial structure of $\mathcal{S}_{+}({G})$.